OpenStudy (anonymous):

Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 2x+6y=-8 5x-3y=88

OpenStudy (anonymous):

( 14, - 6) is what i got

OpenStudy (ajprincess):

If u multiply the second equation what do u get? @Ro_05

OpenStudy (anonymous):

multiply 2 to the second equation that way you have 10x - 6y = 176 , then add the two equations together.

OpenStudy (ajprincess):

@Poetrylover what u got is right. it would be really better if u show the work.:)

OpenStudy (anonymous):

sorry 2x +6y = -8 2(5x - 3y) = 2(88) 2x+6y=-8 10x-6y=176 12x=168 /12 /12 x=14 plug in x 2(14) +6y=-8 28+6y=-8 -28 -28 6y=-36 /6 /6 y=-6 (14, -6)

OpenStudy (ajprincess):

that's k:)